the measure of each base angle of an isosceles 21 more than the measure of the vertex angle. find the measure of each angle of the triangle.
Answer by rapaljer(4671)

Let x = vertex angle
x+21 = each base angle

x+x+21+x+21=180
3x+42=180
3x=138
x= 46 degrees vertex angle
x+21 =67 degrees base angles

Dr. Rapalje